OpenStudy (anonymous):

a ball rolls down an incline and is projected from a height 1.22m at a 40 degree angle, where is the location where the ball will land? Ok so Im thinking, you can try to calculate the Potential energy of the ball at h=1.22m and then relate that somehow to solve for displacement using kinematics equations for 2D motion?

OpenStudy (anonymous):

can you draw it ?

OpenStudy (anonymous):

the drawing is not so great! the radius is the line from the center (approx)

OpenStudy (anonymous):

decided to attach pic

OpenStudy (anonymous):

how about the velocity of the ball ?

OpenStudy (anonymous):

use mechanic energy \[Ep _{0}+Ek _{0}=Ep _{1}+Ek _{1}\] to find the velocity

OpenStudy (anonymous):

v = root(6/5gh) = root (6/5 (9.8 * 1) = 3.42 m/s

OpenStudy (anonymous):

I mean the values of Vo, Vt, and h on that picture

OpenStudy (anonymous):

height is 1m, Vo=O Vt=?

OpenStudy (anonymous):

honestly, im more concerned about the concept than exact values.. so if i find velocity... then do i use translational kinematics to solve for the displacement?

OpenStudy (anonymous):

yes, use parabolic